# The Burgers transform: from holomorphic functions to rigid elliptic structures

### v1

@@ Abstract.
We introduce the Burgers transform $\mathcal{B}$, a nonlinear bijection between holomorphic functions $f\colon U\to\mathbb{C}_{+}$ and rigid variable elliptic structures on the plane, defined implicitly by $\lambda=f(y-\lambda x)$. The output automatically satisfies the conservative complex Burgers equation $\lambda_{x}+\lambda\lambda_{y}=0$.
Our main result is that holomorphicity of the seed $f$ is necessary, not merely sufficient, for rigidity: any $C^{1}$ function whose implicit solution satisfies $\lambda_{x}+\lambda\lambda_{y}=0$ must be holomorphic. This closes a gap in the existing literature and identifies $\operatorname{Hol}(U,\mathbb{C}_{+})$ as the maximal seed space compatible with rigidity. The obstruction formula $H|_{x=0}=2i\,(\operatorname{Im}f)\,f_{\bar{w}}$ quantifies the cost of non-holomorphicity at the level of the initial data.
The transform establishes a hierarchy we call $f$-analytics: standard complex analysis ($f=i$), $p(x)$-analytics ($f\colon U\to i\mathbb{R}_{+}$), and the full rigid class ($f\colon U\to\mathbb{C}_{+}$ holomorphic), in which rigidity—absent from the real diameter of $\mathbb{D}$—emerges as a genuinely complex phenomenon. We characterise the domain of $\mathcal{B}$ through shock formation, its interaction with affine automorphisms of $\mathbb{C}_{+}$, and the infinitesimal the propagator $\mathcal{P}_{f}=D\mathcal{B}_{f}$ satisfies a Jacobian-twisted multiplicativity that deforms the seed algebra by the density of characteristics. Four worked examples—affine, exponential, inverse, and trigonometric seeds—show that the complexity class of a seed and that of the resulting structure are generically unrelated.
@@ 1. Introduction
    A characterisation of $f$-analytics as the maximal seed class compatible with rigidity, with the passage from $p(x)$-analytics governed by the necessity of holomorphicity (Theorem 5.2);
@@
    A characterisation of the domain $\Omega_{f}$ in terms of shock formation (Theorem 4.1);
@@
This paper is a companion to [1], where variable elliptic structures are developed in full. Section 14.4 of that work identifies the implicit equation $\lambda=f(y-\lambda x)$ as the general solution of conservative Burgers, but states the result for $C^{1}$ functions. Theorem 3.1 below corrects this: the rigidity condition forces $f$ to be holomorphic. The Burgers transform was not named or studied as an independent object in [1]; the present paper fills this gap.
@@ 3. Holomorphicity is necessary for rigidity
Let $U\subseteq\mathbb{C}$ be open and connected, and let $f\colon U\to\mathbb{C}_{+}$ be $C^{1}$ in the real sense. If the implicit solution $\lambda=f(y-\lambda x)$ satisfies $\lambda_{x}+\lambda\,\lambda_{y}=0$ on an open set $\Omega\subseteq\mathbb{R}^{2}$ containing a segment of $\{0\}\times\mathbb{R}$, then $f$ is holomorphic on $U$.
@@
For $x\neq 0$, the argument $w=y-\lambda(x,y)\,x$ satisfies $\operatorname{Im}(w)=-x\,\operatorname{Im}\lambda\neq 0$, so $w$ takes genuinely complex values. As $(x,y)$ varies over a neighbourhood of $(0,y_{0})$, the Jacobian of the map $(x,y)\mapsto w$ at $x=0$ has $\partial w/\partial y=1\neq 0$, so by the implicit function theorem the image contains an open subset of $\mathbb{C}$. Hence $f_{\bar{w}}=0$ on an open set in $U$, and since $U$ is connected and $f$ is $C^{1}$, the identity theorem for distributions satisfying $f_{\bar{w}}=0$ gives that $f$ is holomorphic on all of $U$. ∎
@@ 4. The domain of the Burgers transform
Theorem 4.1 (Domain characterisation).
@@ 6.2. Interaction with the affine group
For inversive Möbius transformations $\varphi(\lambda)=(a\lambda+b)/(c\lambda+d)$ with $c\neq 0$, no analogous equivariance holds, even with a compensating coordinate change on the plane. The reason is that $\varphi$ acts nonlinearly on the characteristic speed, and no affine (or even polynomial) reparametrisation of $(x,y)$ can absorb this. The affine subgroup is maximal for equivariance.
@@ 6.3. The residual budget on the Poincaré disk
The Beltrami–Burgers transform inherits the Poincaré geometry of $\mathbb{D}$. The Poincaré residual
@@
Proposition 6.8 (Poincaré residual).
For $f\colon U\to\mathbb{C}_{+}$ of class $C^{1}$, the Poincaré residual of $\mu=\mathcal{C}(\lambda)$ at $x=0$ satisfies
@@ 7.2. Rigid structures as irreducibles
A rigid structure $\mathcal{B}[f]$ is irreducible if there is no factorisation $f=g\cdot h$ with $g,h\in\operatorname{Hol}(U,\mathbb{C}_{+})$ and neither $g$ nor $h$ constant.
@@
Factorisation Decomposition of $f$ in $\operatorname{Hol}(U,\mathbb{C}_{+})$
Unique factorisation? Open
@@ 9.2. The return of holomorphic functions
The theory of variable elliptic structures [1] was designed, in part, to generalise classical complex analysis: the constant imaginary unit $i$ is replaced by a variable $i(x,y)$, and the Cauchy–Riemann equation becomes a family of operators parametrised by the spectral parameter $\lambda(x,y)\in\mathbb{C}_{+}$. In this generalisation, classical holomorphic functions appear as the degenerate case $\lambda\equiv i$, $\mu\equiv 0$—a single point in an infinite-dimensional moduli space. From the perspective of the analysis layer (Vekua equations, Beltrami coefficients, transport dynamics), holomorphic functions are the trivial case, the base point from which all nontrivial structures depart.
@@
    Generators of the rigid core. Not all elliptic structures are rigid, and not all elliptic structures admit a complete function theory. The rigid structures—those with vanishing intrinsic obstruction G=0G=0—are the ones for which the Leibniz rule, the Cauchy–Pompeiu representation, power series, and the similarity principle all hold [1, Chapters 5–9]. The Burgers transform shows that this rigid core is in bijection with $\operatorname{Hol}(U,\mathbb{C}_{+})$. Holomorphic functions are not merely one example of a rigid structure; they are the complete parametrisation of all rigid structures.
The narrative arc is worth stating explicitly, as it illustrates a pattern that may recur in other generalisations of classical analysis:
@@
Generalisation Degenerate special case ($f=i$) $\lambda_{x}+\lambda\lambda_{y}=H$
@@
The generalisation does not eliminate holomorphic functions; it reveals their structural role. In the constant theory, holomorphic functions do everything, and their special nature is invisible—there is nothing to compare them to. In the variable theory, they initially appear to have been superseded. But the rigidity analysis shows that they were never absent: they were operating at a deeper level, as the seeds from which all coherent variable structures grow.
@@ 10.4. The Cauchy kernel
This example also addresses the practitioner’s question of how to translate a classical holomorphic function into the Burgers framework. The naive choice $1/\xi$ maps $\mathbb{R}$ to $\mathbb{R}$ and carries no ellipticity. The regularisation $1/(\xi+i\delta)$ rotates the pole into the lower half-plane and produces a seed with $\operatorname{Im}f>0$ on $\mathbb{R}$.
@@ 11. Open problems
    Unique rigidization. If $\Phi_{1}$ and $\Phi_{2}$ both rigidize the same elliptic structure, must they be related by a rigid automorphism? This is the analogue of unique factorisation.

### v2

@@ Abstract.
We introduce the Burgers transform $\mathcal{B}$, a nonlinear bijection between holomorphic functions $f\colon U\to\mathbb{C}_{+}$, $U\cap\mathbb{R}\supseteq I$ an interval, and rigid variable elliptic structures on the plane, defined implicitly by $\lambda=f(y-\lambda x)$. The output automatically satisfies the conservative complex Burgers equation $\lambda_{x}+\lambda\lambda_{y}=0$.
Our main result is that holomorphicity of the seed $f$ is necessary, not merely sufficient, for rigidity: any $C^{1}$ function whose implicit solution satisfies $\lambda_{x}+\lambda\lambda_{y}=0$ must be holomorphic. This closes a gap in the existing literature and identifies $\operatorname{Hol}(U,\mathbb{C}_{+})$ as the maximal seed space compatible with rigidity. The obstruction formula $H|_{x=0}=2i\,(\operatorname{Im}f)\,f_{\bar{w}}$ quantifies the cost of non-holomorphicity at the level of the initial data. We characterise the domain of $\mathcal{B}$ through shock formation, its interaction with affine automorphisms of $\mathbb{C}_{+}$, and the infinitesimal the propagator $\mathcal{P}_{f}=D\mathcal{B}_{f}$ satisfies a Jacobian-twisted multiplicativity that deforms the seed algebra by the density of characteristics. Four worked examples—affine, exponential, inverse, and trigonometric seeds—show that the complexity class of a seed and that of the resulting structure are generically unrelated.
@@ 1. Introduction
    A characterization of $f$-analytics as the maximal seed class compatible with rigidity, with the passage from $p(x)$-analytics governed by the necessity of holomorphicity (Theorem 5.2);
@@
    A characterization of the domain $\Omega_{f}$ in terms of shock formation (Theorem 4.1);
@@
This paper is a companion to [1], where the fundamental results for variable elliptic structures are developed.
@@ 3. Holomorphicity is necessary for rigidity
Let $U\subseteq\mathbb{C}$ be an open neighbourhood of an interval $I\subset\mathbb{R}$, and let $f\colon U\to\mathbb{C}_{+}$ be $C^{1}$ in the real sense. If the implicit solution $\lambda=f(y-\lambda x)$ satisfies $\lambda_{x}+\lambda\,\lambda_{y}=0$ on an open set $\Omega\subseteq\mathbb{R}^{2}$ containing $\{0\}\times I$, then $f$ is holomorphic on $U$.
@@
For $x\neq 0$, the argument $w=y-\lambda(x,y)\,x$ satisfies $\operatorname{Im}(w)=-x\,\operatorname{Im}\lambda\neq 0$, so $w$ takes genuinely complex values. As $(x,y)$ varies over a neighbourhood of $(0,y_{0})$, the Jacobian of the map $(x,y)\mapsto w$ at $x=0$ is
$\frac{\partial w}{\partial x}\bigg|_{x=0}=-f(y_{0}),\qquad\frac{\partial w}{\partial y}\bigg|_{x=0}=1.$
Since $\operatorname{Im}f(y_{0})>0$, these two vectors are $\mathbb{R}$-linearly independent, so the map $(x,y)\mapsto w$ has rank $2$ at every point $(0,y_{0})$. By the implicit function theorem, its image contains an open neighbourhood of $y_{0}$ in $\mathbb{C}$. As $y_{0}$ ranges over $I$, the union of these neighbourhoods is an open set $V\subseteq U$ containing $I$, on which $f_{\bar{w}}=0$. Since $U$ is itself a neighbourhood of $I$, we may shrink $U$ to $V$ if necessary, obtaining $f_{\bar{w}}=0$ on all of $U$, i.e. $f$ is holomorphic on $U$. ∎
Remark 3.2 (Pointwise holomorphicity).
The obstruction formula (6) is algebraic, not differential: it holds at each point of $\Omega$, not merely on open subsets. In particular, if $\lambda=f(y-\lambda x)$ and $H=\lambda_{x}+\lambda\lambda_{y}=0$ at a single point $(x_{0},y_{0})\in\Omega$ where $\operatorname{Im}\lambda>0$ and $1+f_{w}\,x_{0}\neq 0$, then $f_{\bar{w}}(w_{0})=0$ at $w_{0}=y_{0}-\lambda(x_{0},y_{0})\,x_{0}$. Holomorphicity of $f$ can therefore be read off point by point, wherever the characteristics reach.
Remark 3.3 (The role of the swept region).
The hypothesis that $U$ is a neighbourhood of a real interval (rather than an arbitrary connected open set) ensures that the characteristic sweep $(x,y)\mapsto y-\lambda(x,y)\,x$ covers all of $U$ for $|x|$ small. For general $U$ extending deep into $\mathbb{C}$, the swept region $W=\{y-\lambda(x,y)\,x:(x,y)\in\Omega\}$ may be a proper subset of $U$: reaching a target $w_{0}$ with large $|\operatorname{Im}w_{0}|$ requires $|x_{0}|=|\operatorname{Im}w_{0}|/\operatorname{Im}f(w_{0})$, which may exceed the shock threshold $|1/f^{\prime}(w)\,|$. The conclusion of Theorem 3.1 then holds on $W$ (which is always open and contains $U\cap\mathbb{R}$), but not necessarily on all of $U$. In every application in this paper—and, more generally, whenever $U$ is a neighbourhood of the initial data—$W$ covers $U$ and the distinction is moot.
@@ 4. The domain of the Burgers transform
Theorem 4.1 (Domain characterization).
@@ 6.2. Interaction with the affine group
For inversive Möbius transformations $\varphi(\lambda)=(a\lambda+b)/(c\lambda+d)$ with $c\neq 0$, no analogous equivariance holds, even with a compensating coordinate change on the plane. The reason is that $\varphi$ acts nonlinearly on the characteristic speed, and no affine (or even polynomial) reparametrization of $(x,y)$ can absorb this. The affine subgroup is maximal for equivariance.
@@ 6.3. The residual budget on the Poincaré disk
The Beltrami–Burgers transform inherits the Poincaré geometry of $\mathbb{D}$. The Beltrami residual
@@
Proposition 6.8 (Beltrami residual).
For $f\colon U\to\mathbb{C}_{+}$ of class $C^{1}$, the Beltrami residual of $\mu=\mathcal{C}(\lambda)$ at $x=0$ satisfies
@@ 7.2. Rigid structures as irreducibles
A rigid structure $\mathcal{B}[f]$ is irreducible if there is no factorization $f=g\cdot h$ with $g,h\in\operatorname{Hol}(U,\mathbb{C}_{+})$ and neither $g$ nor $h$ constant.
@@
Factorization Decomposition of $f$ in $\operatorname{Hol}(U,\mathbb{C}_{+})$
Unique factorization? Open
@@ 9.2. The return of holomorphic functions
The theory of variable elliptic structures [1] was designed, in part, to generalise classical complex analysis: the constant imaginary unit $i$ is replaced by a variable $i(x,y)$, and the Cauchy–Riemann equation becomes a family of operators parametrised by the spectral parameter $\lambda(x,y)\in\mathbb{C}_{+}$. In this generalization, classical holomorphic functions appear as the degenerate case $\lambda\equiv i$, $\mu\equiv 0$—a single point in an infinite-dimensional moduli space. From the perspective of the analysis layer (Vekua equations, Beltrami coefficients, transport dynamics), holomorphic functions are the trivial case, the base point from which all nontrivial structures depart.
@@
    Generators of the rigid core. Not all elliptic structures are rigid, and not all elliptic structures admit a complete function theory. The rigid structures—those with vanishing intrinsic obstruction G=0G=0—are the ones for which the Leibniz rule, the Cauchy–Pompeiu representation, power series, and the similarity principle all hold [1, Chapters 5–9]. The Burgers transform shows that this rigid core is in bijection with $\operatorname{Hol}(U,\mathbb{C}_{+})$. Holomorphic functions are not merely one example of a rigid structure; they are the complete parametrization of all rigid structures.
The narrative arc is worth stating explicitly, as it illustrates a pattern that may recur in other generalizations of classical analysis:
@@
Generalization Degenerate special case ($f=i$) $\lambda_{x}+\lambda\lambda_{y}=H$
@@
The generalzation does not eliminate holomorphic functions; it reveals their structural role. In the constant theory, holomorphic functions do everything, and their special nature is invisible—there is nothing to compare them to. In the variable theory, they initially appear to have been superseded. But the rigidity analysis shows that they were never absent: they were operating at a deeper level, as the seeds from which all coherent variable structures grow.
@@ 10.4. The Cauchy kernel
This example also addresses the practitioner’s question of how to translate a classical holomorphic function into the Burgers framework. The naive choice $1/\xi$ maps $\mathbb{R}$ to $\mathbb{R}$ and carries no ellipticity. The regularization $1/(\xi+i\delta)$ rotates the pole into the lower half-plane and produces a seed with $\operatorname{Im}f>0$ on $\mathbb{R}$.
@@ 11. Open problems
    Unique rigidization. If $\Phi_{1}$ and $\Phi_{2}$ both rigidize the same elliptic structure, must they be related by a rigid automorphism? This is the analogue of unique factorization.